Suppose that 10 years ago you bought a home for $130,000, paying 10% as a down payment, and financing the rest at 7% interest for 30 years.


Your existing mortgage (the one you got 10 years ago)


How much money did you pay as your down payment?

How much money was your existing mortgage (loan) for?

What is your current monthly payment on your existing mortgage?

How much total interest will you pay over the life of the existing loan?


1
Expert's answer
2021-03-24T12:01:25-0400

Price of home = $130,000

Less downpayment @ 10% = $130,000*10%

=13,000

Home loan = Price of Home - Down payment

"=130,000-13,000\n=117,000"

$117,000

Interest @ 7% for 30 years

"EMI = [P * R *(1+R)^N]\/[(1+R)^N-1]"

Where P : Loan Amount

R: interest rate

N : number of Monthly Installments

EMI : Monthly installments

"R= 7\/(12 *100) \n =0.0058"

"EMI = [117000 * 0.0058 * (1+0.0058)^360]\/[(1+0.0058)^360-1]"

"= [678.60 * 8.02]\/[7.02]"

"= 775.26"

= $775.26

Total Amount Payed

"=775.26*12*30\n=334,912.32"

Total Interest Payed = Total Amount Payed - Loan Amount

"=334,912.32 - 117,000"

"=217,912.32"


$217,912.32